For cost plus incentive fee contract, calculate the final payment made to the supplier,
Target cost = $100,000
Target profit = $12,000
Maximum fee = $14,000
Minimum fee = $9,000
Sharing below target (customer/supplier) = 80/20
Sharing above target (customer/supplier) = 70/30

1) How much will the contractor be reimbursed if the cost of performing the work is $85,000?
2) How much will the contractor be reimbursed if the cost of performing the work is $120,000?
Answer #1

1.

Final Fee = Target Fee + (Target Cost - Actual Cost) * Contractor Share = 12000 + (100000-85000)*20% = 12000 + 3000 = $15000

This is more than maximum fee, hence the contractor will get maximum permissible fee i.e. $14000.

2. Final Fee = Target Fee + (Target Cost - Actual Cost) * Contractor Share= 12000 + (100000-120000)*30% = $6000

This is less than minimum fee, hence the contractor will get minimum fee i.e. $9000.